# booktabs, siunitx

### v1

@@ 3.4 Primary GWs and constraints from $\Delta N_{\mathrm{eff}}$
which is essentially the same as the equation that describes the Fourier modes of the tensor perturbations at the second order [cf. Eq. (3.1)], but without a source. Consider a slow roll inflationary scenario driven by a single, canonical scalar field. In such a case, it is well known that the spectrum of primary GWs produced at the end of inflation proves to be nearly scale invariant, with an amplitude that is determined by the Hubble scale during inflation. The perturbations that are generated during inflation need to be evolved through the epochs of reheating and radiation domination, in order to arrive at the spectral energy density of primary GWs today (in this context, see, for example, Refs. [200, 201, 202, 203, 204, 205, 206, 207, 19, 208, 209, 210, 211, 212, 213, 214, 215, 216]). The strength and shape of the spectral density of primary GWs today are sensitive to the energy scale of inflation as well as the history of the background evolution after inflation.
For simplicity, let us consider inflation to be of the de Sitter form characterized by the constant Hubble parameter, say, $H_{{}_{\mathrm{I}}}$ [^1] A priori, such an assumption may seem inconsistent with the inflationary model involving an epoch of ultra slow roll that is supposed to produce the scalar power spectrum (2.8). However, it is known that the de Sitter approximation for the scale factor works well during the epoch of ultra slow roll. Moreover, the departure from slow roll only leads to a step in an otherwise scale-invariant tensor power spectrum (in this regard, see, for instance, Ref. [161], Fig. 3).. In such a situation, as is well known, the scale of inflation can expressed in terms of the tensor-to-scalar ratio $r$ and the scalar amplitude $A_{{}_{\mathrm{S}}}$ as follows: $H_{{}_{\mathrm{I}}}=\pi\,M_{{}_{\mathrm{Pl}}}\,\sqrt{r\,A_{{}_{\mathrm{S}}}/2}$. Let us first focus on wave numbers that are significantly larger than the wave numbers associated with the CMB scales and which re-enter the Hubble radius during the epoch of radiation domination. After they enter the Hubble radius, the energy density associated with the GWs behave in the same manner as the energy density of radiation. Given a scale-invariant spectrum of primary GWs at the end of inflation, the dimensionless spectral energy density of primary GWs today, over wave numbers that re-enter the Hubble radius during radiation domination (i.e. over $k<k_{\mathrm{re}}$) proves to be scale invariant as well, with an amplitude that is given by
@@
If there arises a phase of reheating between inflation and the epoch of radiation domination, then, evidently, the wave numbers over the range $k_{\mathrm{re}}<k<k_{\mathrm{e}}$ will re-enter the Hubble radius during the epoch of reheating. Needless to add, the wave number $k_{\mathrm{re}}$ depends on the EoS $w_{\mathrm{re}}$ and the reheating temperature $T_{\mathrm{re}}$. One finds that, over the wave numbers $k_{\mathrm{re}}<k<k_{\mathrm{e}}$, the spectral density of primary GWs today can be expressed as (see, for instance, Ref. [214] for detailed calculation)
@@
$\mu(w_{\mathrm{re}})=(1+3\,w_{\mathrm{re}})^{4/(1+3\,w_{\mathrm{re}})}\,\Gamma^{2}\left(\frac{5+3\,w_{\mathrm{re}}}{2+6\,w_{\mathrm{re}}}\right),\quad n_{w_{\mathrm{re}}}=\frac{2\,(1-3\,w_{\mathrm{re}})}{1+3\,w_{\mathrm{re}}}.$ (3.32)
@@
In Fig. 2, we have plotted the dimensionless spectral density of primary GWs today corresponding to $r=0.036$, $T_{\mathrm{re}}=50\,\mathrm{MeV}$ and values of $w_{\mathrm{re}}$ for which we had plotted the spectral density of secondary GWs generated due to excess scalar power on small scales. We have plotted the primary spectral densities until the frequencies that correspond to $k_{\mathrm{e}}$ for each of the parameters, which is determined by the relation (3.36).
Table 3: We have listed the maximum allowed value of the tensor-to-scalar scalar $r$ that is consistent with the reheating temperature of $T_{\mathrm{re}}=50\,\mathrm{MeV}$, the best-fit values of $w_{\mathrm{re}}$ arrived at on comparison with the NANOGrav 15-year data [cf. Tab. 3.3], and the constraint $\Delta N_{\mathrm{eff}}<0.284$ from Planck [9]. We have listed the values of $r$ for the four models we discussed earlier, viz. R4pF, R3pF, R3pB and R2pB.
@@ 4 Conclusions and discussion
  • Thirdly, in our analysis, we also accounted for the impact of the uncertainties in the collapse condition, in particular, the value of the critical density $\delta_{\mathrm{c}}$, on the formation of PBHs. As one can intuitively predict, assuming a lower value of $\delta_{\mathrm{c}}$ can lead to an acute overproduction of PBHs, whereas a higher value of $\delta_{\mathrm{c}}$ can significantly suppress the number of PBHs formed, thereby allowing a larger volume of the parameter space to be compatible with the stochastic GW background suggested by the NANOGrav 15-year data. This is clearly reflected in the Bayesian factor (listed in Tab. 3.3) for the model R2pB. Note that the Bayesian factor is considerably higher when $\delta_{\mathrm{c}}=1.5\,\delta_{\mathrm{c}}^{\mathrm{an}}$ than when $\delta_{\mathrm{c}}=0.5\,\delta_{\mathrm{c}}^{\mathrm{an}}$.
@@ Acknowledgements
SM and LS wish to thank the Indian Institute of Technology Madras, Chennai, India, for support through the Exploratory Research Project RF22230527PHRFER008479. NB thanks Rajeev Kumar Jain and Ranjan Laha for the initial collaborations and fruitful discussions on the generation of scalar-induced SGWB during the phase of reheating, specifically in the context of the resonant amplification of the SGWB. MRH wishes to acknowledge support from the Science and Engineering Research Board (SERB), Department of Science and Technology (DST), Government of India (GoI), through the National Postdoctoral Fellowship PDF/2022/002988. DM and LS gratefully acknowledge the support received from SERB, DST, GoI, through the Core Research Grant CRG/2020/003664. LS also wishes to thank the Indo-French Centre for the Promotion of Advanced Research for support of the proposal 6704-4 under the Collaborative Scientific Research Programme.
@@
During the completion of this work, a preprint appeared on the arXiv [132], which also studies the generation of secondary GWs during reheating, which are induced by primordial scalar power spectra that contain either a delta function or a lognormal peak on small scales. The work also compares the spectral density of GWs with the NANOGrav data. In the models R4pF and R3pF, we have also considered primordial scalar spectra with enhanced power on small scales, but in the form of a broken power law. To ensure that the optimal parameters do not lead to the overproduction of PBHs, in our work, we have further carried out runs which simultaneously calculate $f_{{}_{\mathrm{PBH}}}(M)$ and scan the parameter space that is restricted to either $f_{{}_{\mathrm{PBH}}}<1$ (in the model referred to as R3pB) or is consistent with the available mass-dependent constraints on $f_{{}_{\mathrm{PBH}}}(M)$ (in the model R2pB). Moreover, in the models R3pB and R2pB, we have explicitly accounted for the uncertainty in $\delta_{\mathrm{c}}$. Further, in the case of the model R2pB, using the enterprise mode of PTArcade, we have evaluated the Bayesian evidences associated with the optimal scenarios for different values of $\delta_{\mathrm{c}}$. We have also considered the primary GWs generated in the scenario and discussed the impact of the constraints from $\Delta N_{\mathrm{eff}}$ on the tensor-to-scalar ratio or, equivalently, the energy scale of inflation.

### v2

@@ 3.4 Primary GWs and constraints from $\Delta N_{\mathrm{eff}}$
which is essentially the same as the equation that describes the Fourier modes of the tensor perturbations at the second order [cf. Eq. (3.1)], but without a source. Consider a slow roll inflationary scenario driven by a single, canonical scalar field. In such a case, it is well known that the spectrum of primary GWs produced at the end of inflation proves to be nearly scale invariant, with an amplitude that is determined by the Hubble scale during inflation. The perturbations that are generated during inflation need to be evolved through the epochs of reheating and radiation domination, in order to arrive at the spectral energy density of primary GWs today (in this context, see, for example, Refs. [229, 230, 231, 232, 233, 234, 235, 236, 19, 237, 238, 239, 240, 241, 242, 243, 244, 245, 246]). The strength and shape of the spectral density of primary GWs today are sensitive to the energy scale of inflation as well as the history of the background evolution after inflation.
For simplicity, let us consider inflation to be of the de Sitter form characterized by the constant Hubble parameter, say, $H_{{}_{\mathrm{I}}}$ [^2] A priori, such an assumption may seem inconsistent with the inflationary model involving an epoch of ultra slow roll that is supposed to produce the scalar power spectrum (2.8). However, it is known that the de Sitter approximation for the scale factor works well during the epoch of ultra slow roll. Moreover, the departure from slow roll only leads to a step in an otherwise nearly scale-invariant tensor power spectrum (in this regard, see, for instance, Ref. [188], Fig. 3).. In such a situation, as is well known, the scale of inflation can expressed in terms of the tensor-to-scalar ratio $r$ and the scalar amplitude $A_{{}_{\mathrm{S}}}$ as follows: $H_{{}_{\mathrm{I}}}=\pi\,M_{{}_{\mathrm{Pl}}}\,\sqrt{r\,A_{{}_{\mathrm{S}}}/2}$. Let us first focus on wave numbers that are significantly larger than the wave numbers associated with the CMB scales and which re-enter the Hubble radius during the epoch of radiation domination. After they enter the Hubble radius, the energy density associated with the GWs behave in the same manner as the energy density of radiation. Given a scale-invariant spectrum of primary GWs at the end of inflation, the dimensionless spectral energy density of primary GWs today, over wave numbers that re-enter the Hubble radius during radiation domination (i.e. over $k<k_{\mathrm{re}}$) proves to be scale invariant as well, with an amplitude that is given by
@@
If there arises a phase of reheating between inflation and the epoch of radiation domination, then, evidently, the wave numbers over the range $k_{\mathrm{re}}<k<k_{\mathrm{e}}$ will re-enter the Hubble radius during the epoch of reheating. Needless to add, the wave number $k_{\mathrm{re}}$ depends on the EoS parameter $w_{\mathrm{re}}$ and the reheating temperature $T_{\mathrm{re}}$. One finds that, over the wave numbers $k_{\mathrm{re}}<k<k_{\mathrm{e}}$, the spectral energy density of primary GWs today can be expressed as (see, for instance, Ref. [244] for detailed calculation)
@@
$\mu(w_{\mathrm{re}})=(1+3\,w_{\mathrm{re}})^{4/(1+3\,w_{\mathrm{re}})}\,\Gamma^{2}\left(\frac{5+3\,w_{\mathrm{re}}}{2+6\,w_{\mathrm{re}}}\right),\quad n_{w_{\mathrm{re}}}=-\frac{2\,(1-3\,w_{\mathrm{re}})}{1+3\,w_{\mathrm{re}}}.$ (3.32)
@@
In Fig. 2, we have plotted the dimensionless spectral density of primary GWs today corresponding to $r=0.036$, $T_{\mathrm{re}}=50\,\mathrm{MeV}$ and values of $w_{\mathrm{re}}$ for which we had plotted the spectral density of secondary GWs generated due to excess scalar power on small scales. We have plotted the primary spectral energy densities until the frequencies that correspond to $k_{\mathrm{e}}$ for each of the parameters, which is determined by the relation (3.33).
Table 3: We have listed the maximum allowed value of the tensor-to-scalar ratio $r$ that is consistent with the reheating temperature of $T_{\mathrm{re}}=50\,\mathrm{MeV}$, the best-fit values of $w_{\mathrm{re}}$ arrived at on comparison with the NANOGrav 15-year data [cf. Tab. 3.3], and the constraint $\Delta N_{\mathrm{eff}}<0.284$ from Planck [9]. We have listed the values of $r$ for the four models we discussed earlier, viz. R4pF, R3pF, R3pB and R2pB.
@@ 4 Conclusions and discussion
  • Thirdly, in our analysis, we also accounted for the impact of the uncertainties in the collapse condition, in particular, the value of the critical density $\delta_{\mathrm{c}}$, on the formation of PBHs. As one can intuitively predict, assuming a lower value of $\delta_{\mathrm{c}}$ can lead to an acute overproduction of PBHs, whereas a higher value of $\delta_{\mathrm{c}}$ can significantly suppress the number of PBHs formed, thereby allowing a larger volume of the parameter space to be compatible with the SGWB suggested by the NANOGrav 15-year data. This is clearly reflected in the Bayesian factor (listed in Tab. 3.3) for the model R2pB. Note that the Bayesian factor is considerably higher when $\delta_{\mathrm{c}}=1.5\,\delta_{\mathrm{c}}^{\mathrm{an}}$ than when $\delta_{\mathrm{c}}=0.5\,\delta_{\mathrm{c}}^{\mathrm{an}}$.
@@ Acknowledgements
SM and LS wish to thank the Indian Institute of Technology Madras, Chennai, India, for support through the Exploratory Research Project RF22230527PHRFER008479. NB wishes to thank his parents for their selfless support that never ceased during this work or any of NB’s earlier works. During this work, NB was supported by the institute postdoctoral fellowship (IPDF) of the Indian Institute of Technology Madras, Chennai, India, and later by postdoctoral fellowship of International Centre for Theoretical Physics Asia-Pacific (ICTP-AP), Beijing, China. NB also thanks Rajeev Kumar Jain and Ranjan Laha for the initial collaborations and fruitful discussions on the generation of scalar-induced SGWB during the phase of reheating, specifically in the context of the resonant amplification of the SGWB. MRH wishes to acknowledge support from the Science and Engineering Research Board (SERB), Department of Science and Technology (DST), Government of India (GoI), through the National Postdoctoral Fellowship PDF/2022/002988. DM and LS gratefully acknowledge the support received from SERB, DST, GoI, through the Core Research Grant CRG/2020/003664. LS also wishes to thank the Indo-French Centre for the Promotion of Advanced Research for support of the proposal 6704-4 under the Collaborative Scientific Research Programme. We also thank our referee in JCAP for his/her fruitful comments, which helped us improve the discussions in the manuscript. During this work, NB was supported by the institute postdoctoral fellowship (IPDF) of the Indian Institute of Technology Madras, Chennai, India, and later by postdoctoral fellowship of International Centre for Theoretical Physics Asia-Pacific (ICTP-AP), Beijing, China.
@@
During the completion of this work, a preprint appeared on the arXiv [137], which also studies the generation of secondary GWs during reheating, induced by primordial scalar power spectra that contain either a delta function or a lognormal peak on small scales. The work also compares the spectral density of GWs with the NANOGrav data. In the models R4pF and R3pF, we have also considered primordial scalar spectra with enhanced power on small scales, but in the form of a broken power law. To ensure that the optimal parameters do not lead to the overproduction of PBHs, in our work, we have further carried out runs which simultaneously calculate $f_{{}_{\mathrm{PBH}}}(M)$ and scan the parameter space that is restricted to either $f_{{}_{\mathrm{PBH}}}<1$ (in the model referred to as R3pB) or is consistent with the available mass-dependent constraints on $f_{{}_{\mathrm{PBH}}}(M)$ (in the model R2pB). Moreover, in the models R3pB and R2pB, we have explicitly accounted for the uncertainty in $\delta_{\mathrm{c}}$. Further, in the case of the model R2pB, using the enterprise mode of PTArcade, we have evaluated the Bayesian evidences associated with the optimal scenarios for different values of $\delta_{\mathrm{c}}$. We have also considered the primary GWs generated in the scenario and discussed the impact of the constraints from $\Delta N_{\mathrm{eff}}$ on the tensor-to-scalar ratio or, equivalently, the energy scale of inflation.